OpenStudy (anonymous):

evaluate the limit lim x approaches 1 (x^2+x-2)/(x-1)....i am stuck... i factor out the number and got the result to be 2 but its wrong..

OpenStudy (anonymous):

factor x^2 + x - 2 into (x-1)(x+2). Then you cancel the (x-1) from the top with the one on the bottom. Then you have lim x-->1 of x+2 = 3.

OpenStudy (anonymous):

oo i mis calulated...thanks

OpenStudy (anonymous):

Just for prettiness:\[\lim_{x \rightarrow 1} \frac{x^2 + x - 2}{x-1}\]\[=\lim_{x \rightarrow 1} \frac{(x-1)(x+2)}{x-1}\]\[=\lim_{x \rightarrow 1} x + 2\]\[=1 + 2\]\[=3\]

OpenStudy (anonymous):

thanks a lot : )